OpenStudy (anonymous):

To exclusively rent a videoke room, a certain number of people will have to pay equal amounts to a total of 720 php. If there were 3 persons less, each person would then have to contribute 40 php more. How many people were there originally?

OpenStudy (ahsome):

Assuming php is a type of currency and 720 is total cost: You have two unknowns, the number of students and the cost per student: Let x be the number of students and y be the cost per student \[xy=720\] \[x-3=y+40\] Simplify 2nd Question \[x=y+43\] Rearrange the first equation \[x=720/y\] Now you know this: \[720/y=y+43\] Solve from there. If you still need further help, just tell me @raechelvictoria

OpenStudy (ahsome):

That will give you the value of y. Sub that into x to get the number of students

OpenStudy (ahsome):

(Need someone to clarify) @iambatman @Kainui